# The PAU Survey: galaxy stellar population properties estimate with narrow-band data

### v1

@@ 3 Methods
This section describes the SED fitting codes we employ for estimating galaxy stellar population properties from photometric data, namely cigale and Prospector, focusing on their different evaluation approaches for inferring SEDs and the corresponding population properties using stellar population synthesis (SPS). We then describe the method we use to create simulated galaxy populations using Prospector with non-parametric star formation history with independent star formation rates over several time bins. This library contains synthetic PAUS and CFHTLS magnitudes and associated errors as well as ground truth physical properties, which can then be utilized for parameter space optimization and general testing of the SED fitting accuracy with narrow-band photometries.
@@
The Code Investigating GALaxy Emission (cigale), introduced in its newest python implementation by Boquien et al. (2019), is a modular SED code covering emission from X-ray and FUV wavelengths up to the the FIR and radio regime. Once a set of models for the desired contributions to the galactic SED has been chosen and a range of values for the free parameters has been set, cigale computes the grid of these models and evaluates them on the data based on their $\chi^{2}$ value, which is calculated according to
@@
Here, $F_{\lambda_{j}}^{\text{obs}}$ is the observed flux in the $j^{\text{th}}$ band, $\sigma_{j}^{2}$ is the corresponding uncertainty of the observation, and $\alpha_{k,l,m}f_{\lambda_{k}}$ are the model fluxes created from a star formation history $\Psi$, a metal enrichment $\zeta$ and additional contributions from for instance dust or nebulae $\xi$. cigale draws simple stellar population (SSP) spectra from the Bruzual & Charlot BC03 stellar evolution synthesis templates (Bruzual & Charlot, 2003) given an initial mass function (IMF) and a set of discrete metallicity values. These spectra are then combined with a parametric star formation history (SFH), attenuated by a flexible attenuation curve as well as corrected for emission from ionized gas and interstellar dust. cigale takes the maximum likelihood model from the the customized grid with a standard likelihood $\mathcal{L}=\exp(-\chi^{2}/2)$ as the resulting SED with its associated properties. For the BC03 templates, cigale provides low- and high-resolution spectra, with the latter facilitating emission line recovery but with an increased run time. Version 2022.1 of the code is used throughout this project.
@@ 3.2 Prospector
While the approach employed by cigale can be computationally fast for low-dimensional problems, the grid size grows exponentially with the number of fitted parameters, thus resulting in high computational requirements for a complex parameter space. Prospector, on the other hand, finds the best-fit SED with Markov-Chain Monte-Carlo (MCMC) sampling from an initial prior space using the Metropolis-Hastings algorithm implemented within the emcee package (Foreman-Mackey et al., 2013). Prospector draws hyper-parameter samples, creates observed-frame SEDs using the FSPS framework, generates mock photometry from the model SED and evaluates the log-likelihood $\log\mathcal{L}(D|\theta,\alpha)$, where D is the data and $\theta,\alpha$ are the model parameters and nuisance parameters. For our purposes, we employ a modified log-likelihood that consists of two separate likelihood terms, one for the narrow-band and one for the broad-band data:
$\log\mathcal{L}_{c}=0.8\,\log\mathcal{L}_{\text{NB}}+0.2\,\log\mathcal{L}_{\text{BB}},$ (2)
Given the larger number of narrow-bands and their higher spectral resolution, we assign a larger weight to the narrow-band (NB) likelihood with respect to the broad-band one.
@@ 3.3 Prospector-$\alpha$ Simulated Stellar Populations
To calibrate the hyper-parameters of cigale and Prospector and use them on real data, we generate mock photometry using the Prospector code itself. We generate a library of 10 000 synthetic galaxies with $1.00.1<z<$ 1.0 and $i_{\mathrm{AB}}<22.5$. These ranges are dictated by the PAUS limiting magnitude and the coverage of the bulk of the PAUS redshift distribution. We use the Prospector-$\alpha$-model described in Leja et al. (2017), Leja et al. (2019) to create the photometric data and the associated set of physical properties. This setup uses a non-parametric SFH, a two-component Charlot & Fall (2000) curve for modelling of the dust attenuation separated by contributions from stellar birth clouds and the diffuse ISM optical depths (with the latter having a modified wavelength scaling that accounts for the UV bump, see Noll et al. 2009, Kriek & Conroy 2013), a dust emission model from Draine & Li (2007), Draine et al. (2014), and nebular emission using the Cloudy templates (Ferland et al., 2017). We first draw 50 000 samples from the prior space described in Tab. 1 and then we randomly select 10 000 sources having $i_{\mathrm{AB}}<22.5$.
@@
 [^1] Model parameters and their priors from which we sample our input for the photometric simulations using the Prospector-$\alpha$-model. $\log(U)$ is the ionisation parameter, $f_{\mathrm{bol}}^{\mathrm{AGN}}$ the AGN fraction of the bolometric luminosity, $\tau_{\mathrm{AGN}}$ the optical depth of the AGN dust torus, $n$ the slope modifier of the Calzetti attenuation curve (Calzetti, 2001), $U_{\min}$ the minimum dust emission radiation field, $\gamma$ its illuminated fraction, $q_{\mathrm{PAH}}$ the fraction of PAHs, $\hat{\tau}_{\lambda,2}$ the diffuse dust optical depth and $\hat{\tau}_{\lambda,1}$ the birth cloud optical depth in the two-component Charlot & Fall (2000) model. In this table, $\mathcal{U}$ specifies a uniform prior that is bound by $\min$ and $\max$ values, $\mathcal{N}$ designates a Gaussian prior characterized by a mean $\mu$ and a standard deviation $\sigma$ and $c\mathcal{N}$ a clipped Gaussian prior that can be used to truncate a specific parameter to exclude for example sampling from negative values.
The SFH priors for the non-parametric SFH are set over seven age bins distributed according to Prospector-$\alpha$ up to the age of the universe at the respective redshift. Therein, the ratios of SFRs $x_{n,n+1}$ in adjacent bins are modelled by independent priors with a Student’s $t$-distribution
$s_{t}(x_{n,n+1},\nu)=\frac{\Gamma\left(\frac{\nu+1}{2}\right)}{\sqrt{\nu\pi}\,\Gamma{\left(\frac{\nu}{2}\right)}}\left(1+\frac{(x_{n,n+1}/\sigma)^{2}}{\nu}\right)^{-\frac{\nu+1}{2}}$ (3)
@@
For each of the 10 000 prior draws, we generate the galaxy SED with Prospector and we obtain synthetic PAUS and CFHTLS model magnitudes. SED fitting routines compute the best-fitting parameters taking into account the photometric errors. Therefore, to realistically mimic the SED fitting measurement of real data, we assign errors to the galaxy model magnitudes. CFHTLS photometric errors are assigned by modeling the magnitude errors as function of magnitude relation using a Gaussian process regressor. To do this, we first split the CFTHLS COSMOS field $ugriz$ data into thin magnitude bins and compute mean and standard deviations of the magnitude errors in each bin. Then, the resulting data is used to train a Gaussian process regressor. The latter is then used to predict the magnitude errors based on the input model magnitudes. The same approach can be implemented for PAUS filters, by using the Flagship-PAU catalogue from Castander et al. (in prep.), Cabayol et al. (2023), which includes photometry of ~130 000 simulated galaxies in the PAUS bands with $i_{\mathrm{AB}}<23$.
Figure 2: Prior distributions of selected parameters used for the generation of mock spectra and photometries for (a) redshift $z$ using the PAUS redshift distribution (red), (b) logarithm of the stellar mass, (c) the mass-weighted stellar age $\left\langle t_{*}\right\rangle_{m}$, (d) the SFR over the last 100 Myrs, (e) the stellar metallicity $Z_{*}$, (f-g) the ISM and BC optical depths $\hat{\tau}_{\lambda,2}$, $\hat{\tau}_{\lambda,1}$, (h) the mass-metallicity relation adopted from Gallazzi et al. (2005) and (i) a random sample of 10 of the non-parametric SFHs.
@@ 4.1 Parameter Space Selection
Table 2: cigale model parameters and values
@@
 [^2] Modules and parameters selected for the cigale SED fitting. Unspecified parameters of the respective model are fixed at their default values. A list of parameters for each module is available in the appendix of Boquien et al. (2019).
@@
In doing so, we decide to keep the redshift fixed to its true input value. A similar choice is performed also on real data, where the redshifts are fixed to the values quoted in the PAUS photo-$z$ catalogue described in detail in Alarcon et al. (2021). Keeping the redshift fixed allows us to reduce the dimensionality of the problem and remove degeneracies between the redshift and the galaxy physical properties.
In Tab. 2, we list the grid of hyper-parameter values for the cigale SED fitting routine. These values have been obtained by finding the right parameter combination that returns the highest degree of consistency between the recovered properties and the input ones. The hyper-parameters are utilised throughout this work, including the estimation of real galaxy properties. In cigale, we use the BC03 simple stellar population templates and a delayed-$\tau$ shape with optional late star-burst or quench for the SFH (Lee et al., 2010; Madau & Dickinson, 2014), for a schematic overview of this SFH’s parameters see for example Fig. 1 in Manzoni et al. 2021. Furthermore, we employ a Charlot & Fall (2000) dust attenuation curve, a parameterized dust emission templates from Draine & Li (2007) and the Cloudy nebular emission templates. The main differences to the simulated galaxies lie in the stellar templates, MILES (Sánchez-Blázquez et al., 2006) vs BC03 for Prospector and cigale, respectively, and in the use of the delayed-tau SFH instead of the non-parametric one. cigale does not contain the latter option in the same manner, as there is a non-parametric SFH, but without the possibility of constraining SFR ratios between adjacent bins.
@@ 4.2 Results with cigale
The hyperparameter grid from Tab. 2 has been determined as the one providing the highest degree of consistency with the input simulated data. We show in Fig. 3 the Bayesian maximum a posteriori estimates from cigale for six SED properties in relation to the true input values of the Prospector-$\alpha$ mock photometries. The properties shown are the stellar mass $M_{*}$, the mean mass-weighted stellar ages $\langle t_{*}\rangle_{m}$, the SFR over the last 100 Myrs, the stellar metallicity $Z_{*}$ and the optical depths of the two-component dust attenuation model $\hat{\tau}_{\lambda,1}$ and $\hat{\tau}_{\lambda,2}$ for the contributions from birth clouds and the ISM, respectively. Additionally, below each 1:1 relation, we show the residuals $r_{i}=\langle p^{C}_{i}\rangle-\langle p^{\alpha}_{i}\rangle$ and their standard deviations $\sigma^{p}_{i}$ for 10 bins covering the parameter range, where $\langle p^{C}_{i}\rangle$ and $\langle p^{\alpha}_{i}\rangle$ are the mean values in the $i^{\text{th}}$ bin from cigale and the simulated library, respectively. The respective bin means are also shown directly in the plot.
@@
Figure 3: Hexbin plots showing the ground truth from the Prospector-$\alpha$ stellar population simulations on the x-axis versus the MAP value from the cigale $\chi^{2}$ analysis, for stellar mass $M_{*}$, mean mass-weighted stellar ages $\langle t_{*}\rangle_{m}$, SFR over the last 100 Myrs, stellar metallicity $Z_{*}$ and the optical depths of the two-component dust attenuation model $\hat{\tau}_{\lambda,2}$ and $\hat{\tau}_{\lambda,1}$. The red points and errorbars show the median and standard deviation for several bins. Below each panel we show mean the residuals and their standard deviations over 10 bins covering the parameter range.
The input SFRs averaged over the last 100 Myrs, show instead a different behaviour. They are well recovered for input SFRs down to roughly $10^{0}M_{\odot}/yr$, but they depart from the 1:1 relation for lower values of the input SFR. In particular, for a fixed input SFR, cigale tends to overpredict the SFR with respect to the simulations, with residuals up to 2-3 dex for input SFRs smaller than $10^{-2}\,M_{\odot}$ yr-1. The results imply that our cigale setup is not able to reliably recover the range below SFR$<10^{0}\,M_{\odot}$ yr-1.
@@
Due to this, mass-weighted ages show a less consistent behaviour with respect to stellar masses and SFRs. The turn-over seems to be at 1 Gyr: galaxies with input ages smaller than 1 Gyr have systematically over-predicted ages, while input ages larger than 1 Gyr lead cigale to under-predict the mass-weighted age, although with smaller residuals with respect to small input ages. The mean mass-weighted stellar age is heavily dependent on the individual SFR ratios of the Prospector-$\alpha$ SFH continuity bins, while the functional form of the delayed-$\tau$ cannot account for multiple changes on the overall rate of star formation, for instance an initial starburst followed by a later quench once the galaxy passes through the green valley. Nevertheless, we would have expected a better agreement of the recovered ages with respect to the input ones due to the addition of the PAUS data, as for example the 4000 Å break, a strong age indicator, is within our baseline for the majority of the sample galaxies. While the PAUS filters do not cover the break in the rest-frame, it shifts into the available wavelength range at approximately $z=0.13$, meaning that $>95\,\%$ of our simulated sample have narrow-band coverage at the 4000 Å break.
@@
The optical depths $\hat{\tau}_{\lambda,1}$ and $\hat{\tau}_{\lambda,2}$ of the two-component dust model have residuals consistent with zero in the range 0.2-2, while they heavily differ at smaller input values. This is not surprising since age, dust and metallicity are degenerate and one parameter influence the other.

### v2

@@ 3 Methods
This section describes the SED fitting codes we employed for estimating galaxy stellar population properties from photometric data, namely cigale and Prospector, focusing on their different evaluation approaches for inferring SEDs and the corresponding population properties using stellar population synthesis (SPS). We also describe the method we used to create simulated galaxy populations using Prospector with non-parametric star formation history with independent star formation rates over several time bins. This library contains synthetic PAUS and CFHTLS magnitudes and associated errors, as well as ground truth physical properties, which can then be utilized for parameter space optimization and the general testing of the SED fitting accuracy with narrowband photometries.
@@
The Code Investigating GALaxy Emission (cigale), introduced in its newest python implementation by Boquien et al. (2019), is a modular SED code covering emission from X-ray and FUV wavelengths up to the FIR and radio regime. Once a set of models for the desired contributions to the galactic SED has been chosen and a range of values for the free parameters has been set, cigale computes the grid of these models and evaluates them on the data based on their $\chi^{2}$ value, which is calculated according to
@@
Here, $F_{\lambda_{j}}^{\text{obs}}$ is the observed flux in the $j^{\text{th}}$ band, $\sigma_{j}^{2}$ is the corresponding uncertainty of the observation, and $\alpha_{k,l,m}f_{\lambda_{k}}$ are the model fluxes created from a star formation history $\Psi$, a metal enrichment $\zeta$ and additional contributions from for instance dust or nebulae $\xi$. cigale draws simple stellar population (SSP) spectra from the Bruzual & Charlot BC03 stellar evolution synthesis templates (Bruzual & Charlot, 2003) given an initial mass function (IMF) and a set of discrete metallicity values. These spectra are then combined with a parametric star formation history (SFH), attenuated by a flexible attenuation curve as well as corrected for emission from ionized gas and interstellar dust. cigale takes the maximum likelihood model from the customized grid with a standard likelihood $\mathcal{L}=\exp(-\chi^{2}/2)$ as the resulting SED with its associated properties. For the BC03 templates, cigale provides low- and high-resolution spectra, with the latter facilitating emission line recovery but with an increased run time. Version 2022.1 of the code is used throughout this project.
@@ 3.2 Prospector
While the approach employed by cigale can be computationally fast for low-dimensional problems, the grid size grows exponentially with the number of fitted parameters, thus resulting in high computational requirements for a complex parameter space. Prospector, on the other hand, finds the best-fit SED with Markov chain Monte Carlo (MCMC) sampling from an initial prior space using the Metropolis-Hastings algorithm implemented within the emcee package (Foreman-Mackey et al., 2013). Prospector draws hyper-parameter samples, creates observed-frame SEDs using the FSPS framework, generates mock photometry from the model SED and evaluates the log-likelihood $\log\mathcal{L}(D|\theta,\alpha)$, where D is the data and $\theta,\alpha$ are the model parameters and nuisance parameters. For our purposes, we employ a modified log-likelihood that consists of two separate likelihood terms, one for the narrowband and one for the broadband data, as:
$\log\mathcal{L}_{c}=0.8\,\log\mathcal{L}_{\text{NB}}+0.2\,\log\mathcal{L}_{\text{BB}}.$ (2)
Given the larger number of narrowbands and their higher spectral resolution, we assigned a greater weight to the narrowband (NB) likelihood with respect to the broadband one.
@@ 3.3 Prospector-$\alpha$ simulated stellar populations
To calibrate the hyper-parameters of cigale and Prospector and use them on real data, we generated mock photometry using the Prospector code itself. We generate a library of 10 000 synthetic galaxies with $1.00.1<z<$ 1.0 and $i_{\mathrm{AB}}<22.5$. These ranges are dictated by the PAUS limiting magnitude and the coverage of the bulk of the PAUS redshift distribution. We used the Prospector-$\alpha$-model described in Leja et al. (2017), Leja et al. (2019) to create the photometric data and the associated set of physical properties. This setup uses a non-parametric SFH, a two-component Charlot & Fall (2000) curve for modeling the dust attenuation separated based on contributions from stellar birth clouds and the diffuse ISM optical depths (with the latter having a modified wavelength scaling that accounts for the UV bump, see Noll et al. 2009, Kriek & Conroy 2013), and a dust emission model from Draine & Li (2007) and Draine et al. (2014), along with nebular emission using the Cloudy templates (Ferland et al., 2017). We first drew 50 000 samples from the prior space described in Table 1 and then we randomly selected 10 000 sources having $i_{\mathrm{AB}}<22.5$.
@@
 [^1] Parameters and their priors from which we sample our input for the photometric simulations using the Prospector-$\alpha$-model: $\log(U)$ is the ionization parameter, $f_{\mathrm{bol}}^{\mathrm{AGN}}$ is the AGN fraction of the bolometric luminosity, $\tau_{\mathrm{AGN}}$ is the optical depth of the AGN dust torus, $n$ is the slope modifier of the Calzetti attenuation curve (Calzetti, 2001), $U_{\min}$ is the minimum dust emission radiation field, $\gamma$ is its illuminated fraction, $q_{\mathrm{PAH}}$ is the fraction of PAHs, $\hat{\tau}_{\lambda,2}$ is the diffuse dust optical depth, and $\hat{\tau}_{\lambda,1}$ is the birth cloud optical depth in the two-component model from Charlot & Fall (2000) . In this table, $\mathcal{U}$ specifies a uniform prior that is bound by $\min$ and $\max$ values, $\mathcal{N}$ designates a Gaussian prior characterized by a mean $\mu$ and a standard deviation, $\sigma;$ finally, $c\mathcal{N}$ is a clipped Gaussian prior that can be used to truncate a specific parameter to exclude, for example, sampling from negative values.
The SFH priors for the non-parametric SFH are set over seven age bins distributed according to Prospector-$\alpha$ up to the age of the universe at the respective redshift. Therein, the ratios of SFRs $x_{n,n+1}$ in adjacent bins are modeled by independent priors with a Student’s $t$-distribution:
$s_{t}(x_{n,n+1},\nu)=\frac{\Gamma\left(\frac{\nu+1}{2}\right)}{\sqrt{\nu\pi}\,\Gamma{\left(\frac{\nu}{2}\right)}}\left(1+\frac{(x_{n,n+1}/\sigma)^{2}}{\nu}\right)^{-\frac{\nu+1}{2}},$ (3)
@@
For each of the 10 000 prior draws, we generated the galaxy SED with Prospector and we obtain synthetic PAUS and CFHTLS model magnitudes. The SED fitting routines compute the best-fitting parameters taking into account the photometric errors. Therefore, to realistically mimic the SED fitting measurement of real data, we assigned errors to the galaxy model magnitudes. The CFHTLS photometric errors are assigned by modeling the magnitude errors as function of magnitude relation using a Gaussian process regressor. To do this, we first split the CFTHLS COSMOS field $ugriz$ data into thin magnitude bins and compute mean and standard deviations of the magnitude errors in each bin. Then, the resulting data was used to train a Gaussian process regressor. The latter was then used to predict the magnitude errors based on the input model magnitudes. The same approach can be implemented for PAUS filters, by using the Flagship-PAU catalogue from Castander et al. (in prep.), Cabayol et al. (2023), which includes photometry of ~130 000 simulated galaxies in the PAUS bands with $i_{\mathrm{AB}}<23$.
Figure 2: Prior distributions of selected parameters used for the generation of mock spectra and photometries for (a) redshift $z$ using the PAUS redshift distribution (red); (b) logarithm of the stellar mass; (c) mass-weighted stellar age $\left\langle t_{*}\right\rangle_{m}$; (d) SFR over the last 100 Myrs; (e) stellar metallicity $Z_{*}$; (f-g) the ISM and BC optical depths, $\hat{\tau}_{\lambda,2}$, $\hat{\tau}_{\lambda,1}$; (h) mass-metallicity relation adopted from Gallazzi et al. (2005), and (i) random sample of ten of the non-parametric SFHs.
@@ 4.1 Parameter Space Selection
Table 2: cigale parameter space
@@
 [^2] Overview of the module and parameter selection for cigale. Unspecified parameters of the respective model are fixed at their default values. A list of parameters for each module is available in the appendix of Boquien et al. (2019).
@@
In doing so, we decided to keep the redshift fixed to its true input value. A similar choice was performed also on real data, where the redshifts are fixed to the values quoted in the PAUS photo-$z$ catalogue described in detail in Alarcon et al. (2021). Keeping the redshift fixed allows us to reduce the dimensionality of the problem and remove degeneracies between the redshift and the galaxy physical properties.
In Table 2, we list the grid of hyper-parameter values for the cigale SED fitting routine. These values have been obtained by finding the right parameter combination that returns the highest degree of consistency between the recovered properties and the input ones. The hyper-parameters are utilised throughout this work, including the estimation of real galaxy properties. In cigale, we used the BC03 simple stellar population templates and a delayed-$\tau$ shape with optional late star-burst or quench for the SFH (Lee et al., 2010; Madau & Dickinson, 2014). For a schematic overview of this SFH’s parameters, we refer to, for example, Fig. 1 in Manzoni et al. 2021. Furthermore, we employed a Charlot & Fall (2000) dust attenuation curve, a parameterized dust emission templates from Draine & Li (2007) and the Cloudy nebular emission templates. The main differences to the simulated galaxies lie in the stellar templates, MILES (Sánchez-Blázquez et al., 2006) vs BC03 for Prospector and cigale, respectively, and in the use of the delayed-tau SFH instead of the non-parametric one. cigale does not contain the latter option in the same manner, as there is a non-parametric SFH, but without the possibility of constraining SFR ratios between adjacent bins.
@@ 4.2 Results with cigale
The hyperparameter grid from Table 2 has been determined as the one providing the highest degree of consistency with the input simulated data. We show in Fig. 3 the Bayesian maximum a posteriori estimates from cigale for six SED properties in relation to the true input values of the Prospector-$\alpha$ mock photometries. The properties shown are the stellar mass, $M_{*}$, the mean mass-weighted stellar ages $\langle t_{*}\rangle_{m}$, the SFR over the last 100 Myrs, the stellar metallicity, $Z_{*}$, and the optical depths of the two-component dust attenuation model $\hat{\tau}_{\lambda,1}$ and $\hat{\tau}_{\lambda,2}$ for the contributions from birth clouds and the ISM, respectively. Additionally, below each 1:1 relation, we show the residuals $r_{i}=\langle p^{C}_{i}\rangle-\langle p^{\alpha}_{i}\rangle$ and their standard deviations $\sigma^{p}_{i}$ for 10 bins covering the parameter range, where $\langle p^{C}_{i}\rangle$ and $\langle p^{\alpha}_{i}\rangle$ are the mean values in the $i^{\text{th}}$ bin from cigale and the simulated library, respectively. The respective bin means are also shown directly in the plot.
@@
Figure 3: Hexbin plots showing the ground truth from the Prospector-$\alpha$ stellar population simulations on the x-axis versus the MAP value from the cigale $\chi^{2}$ analysis, for stellar mass $M_{*}$, mean mass-weighted stellar ages $\langle t_{*}\rangle_{m}$, SFR over the last 100 Myrs, stellar metallicity $Z_{*}$ and the optical depths of the two-component dust attenuation model $\hat{\tau}_{\lambda,2}$ and $\hat{\tau}_{\lambda,1}$. The red points and errorbars show the median and standard deviation for several bins. Below each panel, we show mean the residuals and their standard deviations over ten bins covering the parameter range.
The input SFRs averaged over the last 100 Myrs, instead show a different behavior. They are well recovered for input SFRs down to roughly $10^{0}M_{\odot}/yr$, but they depart from the 1:1 relation for lower values of the input SFR. In particular, for a fixed input SFR, cigale tends to overpredict the SFR with respect to the simulations, with residuals up to 2-3 dex for input SFRs smaller than $10^{-2}\,M_{\odot}$ yr-1. The results imply that our cigale setup is not able to reliably recover the range below SFR$<10^{0}\,M_{\odot}$ yr-1.
@@
As a result, mass-weighted ages show a less consistent behavior with respect to stellar masses and SFRs. The turn-over seems to be at 1 Gyr: galaxies with input ages smaller than 1 Gyr have systematically over-predicted ages, while input ages larger than 1 Gyr lead cigale to under-predict the mass-weighted age, although with smaller residuals with respect to small input ages. The mean mass-weighted stellar age is heavily dependent on the individual SFR ratios of the Prospector-$\alpha$ SFH continuity bins, while the functional form of the delayed-$\tau$ cannot account for multiple changes on the overall rate of star formation, for instance an initial starburst followed by a later quench once the galaxy passes through the green valley. Nevertheless, we would have expected a better agreement of the recovered ages with respect to the input ones due to the addition of the PAUS data, as for example the 4000 Å break, a strong age indicator, is within our baseline for the majority of the sample galaxies. While the PAUS filters do not cover the break in the rest-frame, it shifts into the available wavelength range at approximately $z=0.13$, meaning that $>95\,\%$ of our simulated sample have narrowband coverage at the 4000 Å break.
@@
The optical depths, $\hat{\tau}_{\lambda,1}$ and $\hat{\tau}_{\lambda,2}$, of the two-component dust model have residuals consistent with zero in the range 0.2-2, while they heavily differ at smaller input values. This is not surprising since age, dust and metallicity are degenerate and one parameter influence the other.